40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 9 01:29:26 2017, gautam, Update, LSC, DRMI Nosie Budget v3.0 7x
    Reply  Mon Oct 9 11:55:01 2017, Koji, Update, LSC, DRMI Nosie Budget v3.0 
       Reply  Mon Oct 9 22:18:34 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_dark.pdf
          Reply  Tue Oct 10 22:04:06 2017, rana, Update, LSC, AS55Q Dark Noise 
          Reply  Wed Oct 11 14:42:03 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
             Reply  Wed Oct 11 19:31:32 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
                Reply  Thu Oct 12 01:50:11 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_dark.png
                   Reply  Thu Oct 12 12:17:28 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_AS55_40m.pdf
                      Reply  Fri Oct 13 12:26:12 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_DemodNoises.pdf
                         Reply  Mon Oct 16 16:01:04 2017, gautam, Update, LSC, AS55Q Dark Noise DemodBoardwOscope.pdfDemodBoardwOscope_ASD.pdf
                            Reply  Tue Oct 17 19:31:53 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_TFs.pdfAS55_DemodNoises.pdf
    Reply  Tue Nov 7 17:45:05 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 SRCL_MICH_whitenoise_tf.pdfPRCL_MICH_whitenoise_tf.pdfMICH_aux.pdfC1NB_disp_40m_MICH_NB_2017-10-08.pdf
       Reply  Tue Nov 7 22:56:21 2017, gautam, Update, LSC, DRMI locking recovered DRMI_7Nov20178.png
          Reply  Wed Nov 8 00:28:16 2017, gautam, Update, LSC, Laser intensity coupling measurement attempt PSL_RIN.pdf
       Reply  Wed Nov 8 09:37:45 2017, rana, Update, LSC, DRMI Nosie Budget v3.1 
          Reply  Wed Nov 8 09:59:12 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 
             Reply  Wed Nov 15 01:37:07 2017, gautam, Update, LSC, DRMI low freq. nosie improved C1NB_disp_40m_MICH_NB_2017-11-15.pdfOL_A2L_couplings.pdf
                Reply  Thu Nov 16 00:53:26 2017, gautam, Update, LSC, DRMI noise sub-budgets C1NB_a2l_40m_MICH_NB_2017-11-15.pdfC1NB_aux_40m_MICH_NB_2017-11-15.pdf
Message ID: 13384     Entry time: Tue Oct 17 19:31:53 2017     In reply to: 13382
Author: gautam 
Type: Update 
Category: LSC 
Subject: AS55Q Dark Noise 

[Koji, gautam]

We took a closer look at the AS55 demod board today. The procedure was to just be as thorough as possible, and check the behaviour of the circuit (both Transfer Function and Noise) stage by stage. Checking the transfer function was the key.

During this process, we found that the reason why the Q channels had lower noise than the I channels was because of the gain of the AD829 stage of the circuit was 0dB rather than 4dB (which is what it should be according to the component values used). Specifically, resistor R12, which is supposed to be 1.30kohm, was measured to be 1.03kohmfrown. Replacing this resistor, the transfer functions (see Attachment #1) and noise levels (see Attachment #2) match the expectations from LISO. Some notes:

  1. The daughter board essentially consists of 2 stages
    • OP27 stage, which has a design gain of 16dB ((=316ohm/50ohm) (flat at frequencies <100kHz).
    • AD829 stage, which has a design gain of 4dB (=1.3kohm/768ohm), and is a 2nd order Butterworth LPF with corner @ 1MHz.
    • So the overall gain of the daughter board is 20dB (i.e. x10) at audio frequencies.
  2. The output noise of D040179 is expected to be ~35nV/rtHz at 100Hz, and the measurement (made with inputs soldered together) is consistent with this value.
  3. The measured voltage noise at the input to D040179 (i.e. the output of the minicircuits mixer + SCLF-5 LPF) is ~9nV/rtHz.
  4. The output voltage noise of the demod board with RFPD input terminated then is expected to be the quadrature sum of the noise due to the D040179 electronics (i.e. 40nV/rtHz) and the input noise to the D040179 (i.e. 9nV/rtHz) multiplied by the gain of the daughter board (i.e. x10) == \sqrt{40^2 + 90^2} \approx 98nV/\sqrt{\mathrm{Hz}}.
  5. To calculate the "dark noise" contribution of AS55 to MICH displacement noise, we have to further add the photodiode dark noise contribution: this gets us up to \sqrt{98^2 + 80^2} \approx 130nV/\sqrt{\mathrm{Hz}}. This is consistent with the measurement (see Attachment #2).
  6. Assming the whitened ADC noise level is much below this (should only be ~10nV/rtHz), and given the measured sensing element of 6.2e8 V/m, this means that the dark noise sets a maximum achievable sensitivity of 2e-16m/rtHz.

To figure out what (if anything) is to be done next, we need to first figure out what is the goal. In the end, we care about DARM and not MICH. The optical gain for the former is ~300x the latter, so the dark noise contribution gets scaled by this factor (giving us a number of 7e-19 m/rtHz). There are certainly many noises above that level which have to be handled first. Indeed, looking at the DARM spectrum from DRFPMI lock back in March 2016, it looks like the current 1f DRMI (with coils de-whitened) Michelson sensitivity is within a factor of 2 of DARM in the full lock (albeit with vertex DoFs on 3f signals, and no coil de-whitening). Koji pointed out that we need to consider the photodiode resonant circuit itself too.

TODO: Upload all this onto the DCC

Attachment 1: D040179_TFs.pdf  124 kB  | Hide | Hide all
D040179_TFs.pdf
Attachment 2: AS55_DemodNoises.pdf  231 kB  Uploaded Tue Oct 17 20:36:57 2017  | Hide | Hide all
AS55_DemodNoises.pdf
ELOG V3.1.3-